# Learning to Generate Novel Scientific Directions with Contextualized Literature-based Discovery

### 22 Jan 2024 v4

@@ Abstract
Literature-Based Discovery (LBD) aims to discover new scientific knowledge by mining papers and generating hypotheses. Standard LBD is limited to predicting pairwise relations between discrete concepts (e.g., drug-disease links), and ignores critical contexts like experimental settings (e.g., a specific patient population where a drug is evaluated) and background motivations (e.g., to find drugs without specific side effects). We address these limitations with a novel formulation of contextualized-LBD (C-LBD): generating scientific hypotheses in natural language, while grounding them in a context that controls the hypothesis search space. We present a modeling framework using retrieval of “inspirations” from past scientific papers. Our evaluations reveal that GPT-4 tends to generate ideas with overall low technical depth and novelty, while our inspiration prompting approaches partially mitigate this issue. Our work represents a first step toward building language models that generate new ideas derived from scientific literature.  [^1] Code, data, and resources are publicly available for research purposes: https://github.com/eaglew/clbd.
Learning to Generate Novel Scientific Directions with Contextualized Literature-based Discovery
@@ 1 Introduction
Can machines mine scientific papers and learn to suggest new directions? Nearly four decades have passed since Don Swanson pioneered Literature-Based Discovery (LBD), based on the premise that the literature can be used for generating hypotheses Swanson (1986). LBD focuses on hypothesizing links between previously unstudied pairs of concepts (e.g., new drug-disease links), with a primary application area of drug discovery Henry and McInnes (2017); Sybrandt et al. (2020).
Figure 1: Contextualized Literature-Based Discovery (C-LBD) overview. In C-LBD, the system is given a description of background knowledge and a seed term to generate a new idea. Example from Qin et al. (2022).
While LBD systems have evolved from Swanson’s simple keyword co-occurrence analyses into machine learning approaches Sybrandt et al. (2020), the basic premise remains: predicting pairwise links between concepts from papers. This setting has fundamental drawbacks. Reducing the “language of scientific ideas” Hope et al. (2023) to this simplistic form limits the expressivity of the hypotheses we can hope to generate. Science is done within specific nuanced settings important for qualifying claims mined from literature Sosa and Altman (2022) — e.g., a drug may treat a disease only when administered in a certain way for certain types of patients. Such nuances are difficult to capture with information extraction (IE) systems Luan et al. (2018); Lin et al. (2020); Ye et al. (2022) that require pre-specified schema and are limited in precision and coverage, and are largely ignored in LBD Sosa and Altman (2022). More broadly, LBD does not model contexts that human scientists consider in the ideation process Hope et al. (2023): target application settings, requirements and constraints, motivations and challenges. Finally, the transductive LBD setting, where all concepts are known apriori and only need to be connected to each other, fails to account for the inductive and generative nature of science where new concepts and their recombinations continuously emerge.
We propose Contextual Literature-Based Discovery (C-LBD), a novel setting and modeling framework with a dramatically different and broader approach that addresses these fundamental limitations. Our task is designed in alignment with modern NLP research: it is the first to ground LBD in a natural language context to constrain the generation space, and it also departs from traditional LBD in the output with a sentence generation task. In C-LBD, we are motivated by imagining an AI-based assistant that suggests ideas in natural language, including novel concepts and linkages. As illustrated in Figure 1, the assistant takes as input two contexts: (1) Relevant background, e.g., current problems, motivations and constraints, and (2) a seed term that should be a focus of the generated scientific idea. Given this input, we explore two variants of C-LBD: generating a full sentence describing an idea, and also a more limited objective of generating a salient concept of the idea (see Figure 1). We present a new modeling framework for C-LBD (see Figure 2) that retrieves inspirations from heterogeneous sources derived from scientific literature (e.g., past problems and their solutions, context from a scientific knowledge graph) and uses them to generate new hypotheses. We also introduce an in-context contrastive model which encourages novelty with respect to background sentences.
@@
We perform comprehensive evaluation of language models for generating scientific ideas. We focus on AI/NLP ideas to facilitate analysis by AI researchers themselves, and also demonstrate our approach’s generalization to the biomedical domain. We design extensive evaluation experiments using not only automated metrics but also human annotators with domain expertise to assess relevance, utility, novelty, and technical depth. Our retrieval-augmented hypothesis generation methods substantially improve results; however, analyses show that ideas generated by state-of-the-art models (GPT-4 OpenAI (2023)) fall far behind scientific papers in terms of novelty, depth and utility (§5)—raising fundamental challenges toward building machines that generate scientific directions.
2 Contextual Literature-Based Discovery
Figure 2: Architecture overview. Our models first retrieve heterogeneous inspirations from different sources. We then pass the contextualized input along with retrieved heterogenous neighbors to the generation modules. Finally, an in-context contrastive learning strategy is applied to reduce over-copying.
@@ 2 Contextual Literature-Based Discovery
As a brief recap of Literature-based discovery (LBD) — LBD has been focused on a very specific, narrow type of hypothesis: links between pairs of concepts (often drugs/diseases). The classic formalization of LBD goes back to Swanson (1986) who proposed the “ABC” model where two concepts (terms) A and C are hypothesized as linked if they both co-occur with some intermediate concept B in papers. For example, by analysis of literature using only the co-occurrences of keywords, Swanson discovered a previously unknown link between Raynaud’s disease and fish oil via their shared relationship with blood viscosity, leading to a clinical trial on fish oil for treating the condition DiGiacomo et al. (1989). More recent work has used word vectors Tshitoyan et al. (2019) or link prediction models Wang et al. (2019); Sybrandt et al. (2020) to discover scientific hypotheses as pairwise links between concepts.
@@
To address the limitations discussed in the Introduction, we propose Contextual Literature-Based Discovery (C-LBD). Figure 1 illustrates our setting: given a background description and a seed term to focus on, our aim is to generate idea suggestions. For example, Figure 1 shows a background text that describes problems with “pretrained language models” in the lifelong integration of information sources, including computational costs. The assistant aims to generate an idea for performing “knowledge acquisition” within this context. Importantly, generated ideas should not merely paraphrase the background—the output should be novel with respect to the context. We propose two different formulations of C-LBD: idea sentence generation and idea node prediction. While both tasks utilize the same input format, they diverge in the form of output they target. We describe these task formulations, and how we can obtain large-scale training and evaluation data derived from literature.
Idea-Sentence Generation
Our goal in this setting is to generate a sentence describing an idea, given sentences describing background context and a seed term to guide the generation (see Figure 1). We first use information extraction (IE) to obtain large-scale training data in this form automatically (during training each instance in our automatically-built dataset only contains a seed entity paired with background sentences and the target hypothesis; during evaluation, the target hypothesis is not visible). More formally, we are given a set of documents $\mathcal{D}$, where each document contains the title and abstract of a paper. To obtain background context and meaningful seed terms, we assume each $d\in\mathcal{D}$ is annotated with a graph $G_{d}=(V,E)$ obtained by information extraction systems (§A). Each node $v\in V$ is a cluster of coreferential scientific entity mentions representing a concept, and each directed edge $e_{ij}\in E$ is a relation between two mentions $v_{i}$ and $v_{j}$ (e.g., the [method,used-for,task] relation). Each directed edge $e_{ij}\in E$ is then paired with a context sentence $s_{ij}\in d$—the sentence mentioning $e_{ij}$. Each document $d\in\mathcal{D}$ also comes annotated with a background context $\mathcal{B}$A)). We treat the remaining sentences $s_{ij}$ as target sentences containing new relationships $e_{ij}$. Given a node $v_{i}\in G_{d}$, background context $\mathcal{B}$ from a specific document $d$, and desired target node type $p$, our goal is to generate a full sentence $s_{ij}$ describing a relationship between two entities $v_{i}$ and $v_{j}$, where $v_{j}$ may be a new node not already present in the graph $\mathcal{G}_{d}$. This is illustrated in Figure 1, where $v_{i}$ corresponds to “knowledge acquisition”, $\mathcal{B}$ to the Background Context, and $s_{ij}$ is the Description Sentence which describes a used-for relation in which a method (“function preserved model expansion”) is used for the task of knowledge acquisition.
Idea-Node Prediction
Apart from idea sentence generation, we also formulate hypothesis generation as a link prediction task to more closely resemble the traditional hypothesis generation setting. Unlike standard work in LBD/hypothesis generation, our setting is of contextual prediction, where the context is provided via natural language sentences as in the above. Given the context, we predict new links between existing concepts, and generate new concepts. Our contextual node prediction is formulated using similar notations as above: given a node $v_{i}\in G_{d}$, background context $\mathcal{B}$ from a specific document $d$, and desired target node type $p$ belonging to one of six types (Task, Method, Evaluation Metric, Material, Other Scientific Terms, Generic Terms), the model needs to predict a new node $v_{j}\in\mathcal{G}$ related to it (e.g., in Figure 1, predicting “function preserved model expansion” given the seed+context), where $\mathcal{G}=\cup_{d\in\mathcal{D}}G_{d}$ is a knowledge graph which covers all papers in the corpus $\mathcal{D}$, built as a union of all paper-specific graphs extracted automatically with IE as described above.
3 Models
We now present our models for the new tasks. We present a new module to retrieve heterogeneous inspirations to enrich the contextual input (§3.1). Then, we describe another module to generate a hypothesis (idea) given the enriched context (§3.2). Finally, we introduce a new in-context contrastive augmentation (§3.3). For details on the training setups and hyperparameter tuning for the models described below, see Appendix B.
3.1 Inspiration Retrieval Module
We take broad inspiration from cognitive aspects of innovation Hope et al. (2023): when researchers generate a new idea, they are grounded in a web of existing concepts and papers bearing on the new idea. We aim to enrich the context of each background+seed by retrieving “inspirations”—related pieces of information from different sources that can guide the hypothesis generation.
As illustrated in Figure 2, for a given instance of the C-LBD task, our retrieval augmentation can retrieve different neighbors from three types of graphs (described below) and add them as additional input. Given the seed term $v$ and its corresponding background context $\mathcal{B}$ from document $d$, we construct a string $q$ which is the concatenation of $\mathcal{B}$ and the prompt $\mathcal{P}$ which is based on $v$ and one of two templates: “$v$ is used for $p$” or “$v$ is done by using $p$”. [^2] Other templates we experimented with had similar performance in our experiments, but future work could explore optimal prompting strategies for hypothesis generation. $p$ is the target node type.
For example, in Figure 2, the concatenation is “Knowledge acquisition is done by using Method Context: Current pre-trained language models (plm) are typically trained with static data…”. We use $q$, $v$, and $d$ as queries for each graph respectively. We concatenate each particular type of retrieved neighbor with the input in the following experiments. We now describe the three graphs we leverage: a semantic similarity graph, a knowledge graph, and a citation graph.
Semantic Neighbors
Input-target pairs in the training set can serve as a guiding reference for predicting a new target given an input. To find input-target pairs that are relevant for a specific query $q$, we construct a fully connected graph $\mathcal{G}_{S}$ based on the training set, where each node is a pair of input text $q_{i}$ and target entity $u_{i}$. We define the weight between two nodes $i$ and $j$ as the cosine similarity between $q_{i}$ and $q_{j}$ based on representations from an encoder such as SentenceBERT (Reimers and Gurevych, 2019). Given $q$, we first insert it into $\mathcal{G}_{S}$ and compute the weights of its connected edges. We then retrieve neighbors =$\mathcal{R}=\{R_{1},R_{2},...,R_{k}\}$ from the training set with the largest edge weight, where $k$ is the number of retrieved instances, and $R_{i}$ is a pair of input text $q_{i}$ and target node $u_{i}$. We consider the target node set $\mathcal{U}_{\mathcal{D}}=\{u_{1},u_{2},...,u_{k}\}$ as potentially relevant nodes. For example, in Figure 2, we find the string “informative entities are done by using Method context: in this work, we aim at equipping pre-trained language models with structured knowledge.” as similar to the input and select its target “linked knowledge graph”. We concatenate those entities to form the relevant document context $[u_{1};u_{2};...;u_{k}]$.
KG Neighbors
In addition to using semantic relatedness based on neural text embeddings to find relevant training examples, we also explore enriching the context by linking it to a background KG with useful information on related methods and tasks. In particular, given the seed term $v$, we select one-hop connected neighbors denoted as $\mathcal{N}_{v}=\{n_{1},n_{2},...\}$ from the background KG $\mathcal{G}_{B}$, where the background KG $\mathcal{G}_{B}=\cup_{d\in\mathcal{D}_{\mathcal{T}}}G_{d}$ is the KG which covers all papers in the corpus $\mathcal{D}_{\mathcal{T}}$ prior to a given year $\mathcal{T}$. As an example, in Figure 2, the neighbor nodes of “knowledge acquisition” include “collaborative web text annotation editor”, “image matching”, etc., which we select as additional context. We concatenate those entities to form the KG context $[n_{1};n_{2};...]$.
Citation Neighbors
Another notion of contextual relatedness we explore is via citation graph links, which may capture relevant context missed by textual similarity measures. Given the query document $d$, we consider its cited paper title set $\mathcal{C}_{d}$ as potential candidates. Because the training set only contains paper before year $\mathcal{T}$, we only select papers $\mathcal{C}_{d\mathcal{T}}\subseteq\mathcal{C}_{d}$ prior to year $\mathcal{T}$. We then retrieve the top-$k$ titles with the highest cosine similarity to the query $q$ from the selected cited paper $\mathcal{C}_{d\mathcal{T}}$ based on the representation provided by the encoder. For instance, in Figure 2, the paper ELLE Qin et al. (2022) cites the paper de Masson d'Autume et al. (2019). Therefore, we choose the paper title “episodic memory in lifelong language learning” as additional information. We concatenate those titles as cited paper context $[c_{1};c_{2};...]$ where $c_{i}\in\mathcal{C}_{d\mathcal{T}}$. By retrieving from cited papers, we limit the set of candidates in the semantic graph $\mathcal{G}_{S}$ to a more relevant subset. This can be seen as a stronger assumption on information available to the model— assuming a researcher using the model provides background context in the form of relevant candidate documents from which ideas could be pooled.
3.2 Generation Module
Given retrieved inspirations, a generation module generates the hypothesis. We include several baselines in text generation and node prediction tasks.
Causal Language Modeling
We experiment with recent state-of-the-art LLMs, GPT3.5 davinci-003 Ouyang et al. (2022) and GPT4 gpt-4-0314 checkpoint OpenAI (2023). We first ask the model to generate sentences based on the seed term and the context in the zero-shot (ZS) setting without any in-context examples (GPT3.5ZS, GPT4ZS). We then ask the model to generate sentences or the top-10 node predictions in a few-shot (FS) setting by prompting randomly chosen pairs of input and output from the training set (GPT3.5FS, GPT4FS). Inspired by Liu et al. (2022), we further employ a few-shot setting using semantically similar examples. Instead of random in-context examples, we use the top-$k$ examples from the training set with the highest cosine similarity to the query GPT3.5Retr. This few-shot retrieval setting differs from the semantic retrieval discussed above, in that we provide both the input and output for each instance rather than solely supplying target entities as additional context.
Sequence-to-Sequence Fine Tuning
We fine-tune T5 Raffel et al. (2020); Ribeiro et al. (2021) (more recent models may be used too; see our biomedical experiment §5). T5 has recently been used for KG link prediction: KGT5 Saxena et al. (2022) uses compositional entity representations and autoregressive decoding for link prediction inference to reduce model size while outperforming traditional models. We apply diverse beam search Vijayakumar et al. (2017) for decoding to increase the predictions’ diversity (T5+Div). We also experiment with constrained beam search Anderson et al. (2017) (T5+CBS) to decrease the search space by covering only sequences observed in the bank of entities in the data (training+dev+test).
Dual-Encoder
For node prediction, we explore a dual-encoder Wang et al. (2021) which relies on encoder language models such as BERT Devlin et al. (2019) and has recently been used in link prediction Safavi et al. (2022). Recently, Wang et al. (2022) proposed SimKGC which uses a contrastive learning architecture to improve efficiency; we use SciBERT Beltagy et al. (2019) and CS_RoBERTa Gururangan et al. (2020) (SciBERT and CSRoBERTa).
3.3 In-context Contrastive Augmentation
We observe that the generation models tend to copy phrases from the background context. For example, given the context “…hierarchical tables challenge numerical reasoning by complex hierarchical indexing…”, the model will generate “numerical table reasoning” as the top prediction. For generating suggestions of novel ideas, we may wish to discourage overly copying from the background information. We introduce a new in-context contrastive objective, where negative examples are taken from the text in the input context (e.g., in Figure 1, the in-context negatives are plms, pre-training, etc). Specifically, for the dual-encoder model, we directly add the negatives to the existing InfoNCE loss. For the sequence-to-sequence model, we follow Wang et al. (2023) to compute an InfoNCE loss over the hidden states of the decoder, aiming to maximize the probability of the ground truth against the probabilities of in-context negatives:
Dataset Split Forward Backward Total
Sentence Generation

### 17 Feb 2024 v5

# SciMON[Uncaptioned image]  : Scientific Inspiration Machines Optimized for Novelty

@@ Abstract
We explore and enhance the ability of neural language models to generate novel scientific directions grounded in literature. Work on literature-based hypothesis generation has traditionally focused on binary link prediction—severely limiting the expressivity of hypotheses. This line of work also does not focus on optimizing novelty. We take a dramatic departure with a novel setting in which models use as input background contexts (e.g., problems, experimental settings, goals), and output natural language ideas grounded in literature. We present SciMON, a modeling framework that uses retrieval of “inspirations” from past scientific papers, and explicitly optimizes for novelty by iteratively comparing to prior papers and updating idea suggestions until sufficient novelty is achieved. Comprehensive evaluations reveal that GPT-4 tends to generate ideas with overall low technical depth and novelty, while our methods partially mitigate this issue. Our work represents a first step toward evaluating and developing language models that generate new ideas derived from the scientific literature. [^1] Code, data, and resources are publicly available for research purposes: https://github.com/eaglew/clbd.
SciMON[Uncaptioned image]  : Scientific Inspiration Machines Optimized for Novelty
@@ 1 Introduction
Can machines mine scientific papers and learn to suggest new directions? The idea that information from the literature can be used for automatically generating hypotheses has been around for decades Swanson (1986). To date, the focus has been on a specific setting: hypothesizing links between pairs of concepts (often in drug discovery applications Henry and McInnes (2017), e.g., new drug-disease links), where concepts are obtained from papers or knowledge bases previously derived from papers Sybrandt et al. (2020); Nadkarni et al. (2021).
Figure 1: SciMON takes background context and generates ideas grounded in literature inspirations, optimizing novelty by iteratively comparing to related work.
This common setting has fundamental drawbacks. Reducing the “language of scientific ideas” Hope et al. (2023) to this simplistic form limits the expressivity of the hypotheses we can hope to generate, and does not capture nuanced contexts that scientists consider: target application settings, requirements and constraints, motivations and challenges. In light of the strong progress recently made with large language models (LLMs), in this paper we explore a dramatically different setting: models that take descriptions of problem contexts—and return natural language suggestions of novel scientific directions that are grounded in literature.
We develop a framework named SciMON (Scientific Inspiration Machines with Optimization for Novelty), named after Nobel laureate and AI pioneer Herbert Simon who authored early foundational work on automated scientific discovery Newell and Simon (1956); Simon (1973). We first present an automated data collection methodology that collects examples of past problems and proposed ideas from scientific papers. We then use this data for both fine-tuning and in-context training of LLMs—training them to take problem descriptions and output proposed ideas to address them. We observe that state-of-art LLMs (e.g., GPT-4 OpenAI (2023)) struggle with generating novel scientific ideas, and contribute a new modeling framework for generating hypotheses that makes progress in improving the hypothesis generation ability of LLMs (Figure 1). Given a background problem description, models first dynamically retrieve inspirations from past literature in the form of related problems and their solutions along with contexts from a scientific knowledge graph. These retrieved inspirations serve to ground the generated ideas in existing literature. We then endow models with the ability to iteratively boost the novelty of generated ideas. Given an idea $\mathcal{I}$ generated by the LLM at step $t$, the model compares $\mathcal{I}$ with existing research in the literature; if it finds strongly overlapping research, the model is tasked with updating its idea to be more novel relative to prior work (much like a good researcher would do). We also introduce an in-context contrastive model which encourages novelty with respect to background context.
@@
We perform the first comprehensive evaluation of language models for generating scientific ideas in our new generative, contextual setting. We focus on AI/NLP ideas to facilitate analysis by AI researchers themselves, and also demonstrate generalization to the biomedical domain. We design extensive evaluation experiments using human annotators with domain expertise to assess relevance, utility, novelty, and technical depth. Our methods substantially improve the ability of LLMs in our task; however, analyses show that ideas still fall far behind scientific papers in terms of novelty, depth and utility—raising fundamental challenges toward building models that generate scientific ideas.
2 Background and New Setting
Figure 2: Architecture overview. Our models retrieve inspirations and then pass the background input and retrieved inspirations to an LM-based generation module, which iteratively optimizes novelty. Input from Qin et al. (2022).
We begin with a brief description of related work and background. We then present our novel setting.
@@ 2 Background and New Setting
Literature-based discovery Nearly four decades have passed since Don Swanson pioneered Literature-Based Discovery (LBD), based on the premise that the literature can be used for generating hypotheses Swanson (1986). LBD has been focused on a very specific, narrow type of hypothesis: links between pairs of concepts (often drugs/diseases). The classic formalization of LBD goes back to Swanson (1986) who proposed the “ABC” model where two concepts (terms) A and C are hypothesized as linked if they both co-occur with some intermediate concept B in papers. More recent work has used word vectors Tshitoyan et al. (2019) or link prediction models Wang et al. (2019); Sybrandt et al. (2020) to discover scientific hypotheses as pairwise links between concepts. A tightly related body of research focuses on scientific knowledge graph link prediction Nadkarni et al. (2021), where predicted links may correspond to new hypotheses, and knowledge bases are reflections of existing scientific knowledge in specific domains, derived from literature. A fundamental gap in this line of work is in the lack of approaches for modeling nuanced contexts Sosa and Altman (2022) (e.g., the specific settings in which a drug may be relevant for a disease) for generating ideas in open-ended problem settings with unbounded hypothesis spaces, and for optimizing novelty. Our setting can be viewed as a radical departure addressing the limitations in existing settings.
LLMs for Scientific Innovation Large language models (LLMs) have made remarkable progress in interpreting and producing natural language content and handling knowledge-intensive tasks such as in the medical domain Nori et al. (2023). Very recent work Boiko et al. (2023) has explored the use of LLMs in a robotic chemistry lab setting, planning chemical syntheses of known compounds and executing experiments. Robotic lab settings are inherently limited to narrow sub-areas where such experiments are possible and relevant. Other very recent work Huang et al. (2023) used LLMs to produce code for machine learning tasks such as Kaggle competitions, finding that a GPT-4 agent achieved 0% accuracy on research challenges such as BabyLM Warstadt et al. (2023). GPT-4 has been anecdotally reported as having “strengths less like those of having a human co-author, and more like a mathematician working with a calculator” Carlini (2023). Our goal is to conduct a non-anecdotal evaluation and enhancement of strong LLMs’ ability to generate novel open-ended scientific ideas.
2.1 SciMON Problem Setting
@@
We are motivated by imagining an AI-based assistant that suggests ideas in natural language. The assistant takes as input background context $\mathcal{B}$ consisting of (1) current problems, motivations, experimental settings and constraints, denoted as $\mathcal{M}$; and optionally (2) a seed term $v$ that should be a focus point of the generated idea $\mathcal{I}$. The seed term is motivated by considering a user-provided cue for the model to limit its hypothesis space. Importantly, generated ideas should not merely paraphrase the background—the output should be novel with respect to $\mathcal{B}$ and the broader literature corpus. Figure 2 illustrates the setting, showing a background text that describes problems with “pretrained language models” in the lifelong integration of information sources, including computational costs. The assistant aims to generate an idea for performing “knowledge acquisition” within this context. Given this input, we aim to generate a full sentence describing a novel idea.
Figure 3: We use IE to obtain literature data for our approach: problems/motivations (background) and proposed ideas (target), as well as salient seed terms.
2.2 Automated Training Data Collection
We obtain training data derived from papers with scientific information extraction (IE) models—extracting past examples of background sentences and corresponding ideas (e.g., descriptions of methods used for specific problems in the background sentences), along with salient entities as seed terms. This data is used for training in both in-context learning and fine-tuning setups.
We construct a corpus $\mathcal{D}$ from 67,408 ACL Anthology papers from S2ORC Lo et al. (2020) (we later also conduct an experiment with a biomedical corpus §4.1). Given a title and the corresponding abstract from a document $d$, to select problem/motivation sentences $\mathcal{M}$ we first perform scientific sentence classification Cohan et al. (2019) to classify sentences from the abstract into one of {Background, Method, Objective}, selecting sentences with labels of Background and treating the remaining sentences as target sentences $\mathcal{T}$ which will serve as desired output examples (Figure 3).
For seed term selection, we apply a state-of-the-art scientific IE system Ye et al. (2022) to $\mathcal{T}$ to extract entities corresponding to Task, Method, Evaluation Metric, Material, and relations of the form [method,used-for,task]—mentions of methods and the tasks they are used for, materials used for tasks, etc. We treat the head (e.g., method) or tail (e.g., task) entity as the seed term, and name the other entity (tail/head, respectively) as a target term $t\in\mathcal{T}$. Continuing our example from Figure 2, Figure 3 shows how the seed and target terms (“knowledge acquisition” and “function preserved model expansion”) are extracted from $\mathcal{T}$. During training, each instance contains ($\mathcal{B}$,$\mathcal{T}$) pairs; during evaluation, target information is removed.
We use SciCo Cattan et al. (2021) to obtain coreference links for entity normalization, and use ScispaCy Neumann et al. (2019) to replace abbreviations with a more informative long form. We also collect paper metadata, including the citation network $\mathcal{G}_{c}$. We split our dataset temporally (train/dev/test correspond to papers from years $<$2021 / 2021 / 2022 respectively). For our experiments, we used model checkpoints trained on data preceding 2022, avoiding the risk of data contamination (§6). Table 1 shows data statistics. [^2] More details are in Appendix C.
Quality of IE Preprocessing During preprocessing, we only keep high-confidence outputs from IE models to reduce errors. We observe this removes many of the noisy cases. To validate this, we manually evaluate the precision of each preprocessing step on a random sample of papers and observe that all steps yield high precision (91%-100%) except relation extraction (65%); in total, the rate of instances passing all steps was 79.7%. [^3] See Table 6 in Appendix.
Gold Test Set We create a high-quality, clean test set. We remove test instances where models can trivially use surface-level background information to infer the ground truth to create a more challenging set, selecting instances with low similarity between background and ground truth sentences. We compute the cosine similarity between each instance’s background and corresponding ground truth sentence in the test set and take pairs with similarity $\leq 0.074$, which amounts to the tenth percentile of pairs. We further annotate this subset to create a gold subset. We manually exclude instances with trivial overlap between ground truth and background, remove cases with irrelevant background, and retain only instances where the target relation (from which the seed term is taken) is salient to the target sentence. We also remove test pairs that have unexplained terms in the background. We obtain a total of 194 instances. [^4] Full annotation details are in Appendix C.
Split Forward Backward Total